OpenStudy (anonymous):

what is the solution of the following system x-y=11,-x+y=-11 a. (-3,-4) b. no solutions c. infinitely many solutions d(3,4)

OpenStudy (anonymous):

Are these 2 different equations....or really the same equation? Can you manipulate one of them to make the other?

OpenStudy (anonymous):

i think so....

OpenStudy (anonymous):

Yes. Take either equation. Lets take the 2nd one. Multiply both sides by -1 -1(-x+y) = -11 (-1) x - y = 11..........and we have the 1st equation. How many solutions are there for x - y = 11?

OpenStudy (anonymous):

12 - 1 = 11 13 - 2 = 11 14 - 3 = 11 15 - 4 = 11 16 - 5 = 11 17.5 - 6.5 = 11 and on and on and on and on.............................

OpenStudy (anonymous):

so its c right

OpenStudy (anonymous):

correct. A linear equation like x - y = 11 always has infinite solutions.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem.